One year ago, you invested $2,850. Today, it is worth $3,650. What rate of interest did you earn?

A. 3.56%

B. 14.04%

C. 4.56%

D. 2.34%

E. 28.07%
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

3650=2850*(1+r/100)^1

(3650/2850)=(1+r/100)

(1+r/100)=1.2807

r=1.2807-1

=28.07%(Approx).